OpenStudy (anonymous):

Lisa purchased a hat for $10, a purse for $20, and a book for $23. The sales tax rate was 7.2 percent. What was the total amount she paid?

OpenStudy (anonymous):

First add up the item pricesl: 10 + 20 + 23 = 53 Put that aside for now. 7.2% = 0.072 Find the sales tax of each item 10 * 0.072 = .72 20 * 0.072 = 1.44 23 * 0.072 = 1.66 Total sales tax: .72 + 1.44 + 1.66 = $3.82 Add the Sales tax to the price of the items. 53 + 3.82 = $56.82

OpenStudy (anonymous):

thank you jazy

OpenStudy (anonymous):

welcome. (: